Pat's father is 30 years older than Pat. In 17 years time he will be 5 years less than 3 times Pat's present age. How old is Pat? Let Pat's age be x. So father's age is ...........?
Answer by ewatrrr(24785)

 
Hi
Pat's father is 30 years older than Pat
Let x and (x+30)represent Pat and his father ages NOW
Question states*** in 17 years...
(x+30) + 17 = 3x - 5
Solving for x
x + 47 = 3x - 5
52 = 2x
x = 26, Pat's age NOW. His father is now 56

CHECKING our Answer***
73 = 3*26 - 5 = 78 - 5 = 73